# Learning physical unknowns from hydrodynamic shock and material interface features in ICF capsule implosions

### v1

@@ Abstract
In high energy density physics (HEDP) and inertial confinement fusion (ICF), predictive modeling is complicated by uncertainty in parameters that characterize various aspects of the modeled system, such as those characterizing material properties, equation of state (EOS), opacities, and initial conditions. Typically, however, these parameters are not directly observable. What is observed instead is a time sequence of radiographic projections using X-rays. In this work, we define a set of sparse hydrodynamic features derived from the outgoing shock profile and outer material edge, which can be obtained from radiographic measurements, to directly infer such parameters. Our machine learning (ML)-based methodology involves a pipeline of two architectures, a radiograph-to-features network (R2FNet) and a features-to-parameters network (F2PNet), that are trained independently and later combined to approximate a posterior distribution for the parameters from radiographs. We show that the estimated parameters can be used in a hydrodynamics code to obtain density fields and hydrodynamic shock and outer edge features that are consistent with the data. Finally, we demonstrate that features resulting from an unknown EOS model can be successfully mapped onto parameters of a chosen analytical EOS model, implying that network predictions are learning physics, with a degree of invariance to the underlying choice of EOS model.
@@ 1.1 Inferring physical parameters from radiographic data
Consequently, characterization of the initial conditions responsible for the instability as well as material properties characterizing the growth rates demands new techniques to solve this inverse problem. One such approach to resolve this difficulty has recently been proposed,[59] which utilizes the robust features of the outgoing shock to characterize the growth rates of the instability via a density reconstruction of the radiographic images. The success of this methodology is founded in the fact that the outgoing shock encodes sufficient information to enable machine learning techniques to learn a mapping between a sequence of outgoing shocks and the corresponding density fields. In this work we propose to utilize the outgoing shock as well as outer material edge as robust features to enable parameter estimation and estimation of the initial conditions.
@@ 1.2 A machine learning (ML)-based inverse mapping based on shock and outer edge features
Here we develop a framework to construct machine-learning-based inverse mappings directly from experimental radiographic images to the underlying physical parameters and initial perturbations in ICF settings systems. We assume a governing physical model given by the compressible Euler equations, dropping higher-order effects of radiation for ease of testing and development, analogous to recent work [60], and we assume unknown initial conditions and material properties for the governing Euler equations. As such, we build upon previous work that demonstrates the ability to both learn a sequence of density fields as well as parameters from a sequence of radiographic features to diagnose asymmetries in the drive of ICF capsules, by utilizing an inert gas as a surrogate for the D-T fuel to enable extraction of the outgoing shock [61, 47, 62]. We note that recent work at the National Ignition Facility has been performed using a silicon dopant  [63]. Furthermore, by subsequent design optimization to minimize asymmetries, improved neutron yields can be realized. Our specific problem description is one in which an ICF-like shell is imploded with an initially perturbed surface as depicted in Figure 3. Upon collapse, the generation of a shock forms on axis and subsequently rebounds, interacting with the perturbed surface of the shell and initiates a RM instability. We posit that the outer material edge, outgoing shock, and their evolution encode sufficient information from the instability to identify the underlying simplified Mie-Grüneisenn EOS parameters and structure of the initial shell perturbation and velocity. Moreover, the shock and edge profiles are some of the few robust and identifiable features in dynamic radiographic images of HEDP experiments, which in general are subject to noise and scattering effects, and correspond to a projected areal mass rather than a primary hydrodynamic variable such as density. Thus broadly, we aim to calibrate material models and initial conditions to be consistent with the outgoing shock and edge profile, a problem of data-assimilation [64]. There are many approaches to data assimilation, and here we review some of the prominent techniques, before concluding with our specific machine-learning-based approach.
@@
Consequently, in this work we develop machine learning architectures that directly take in a low-dimensional time-series of radiographic images, extract the outgoing shock and outer material edge profiles, and estimate the corresponding EOS parameters as well as initial shell conditions. The generation of data is discussed in Section 2 and the machine learning (ML)-based parameter estimation pipeline is introduced in Section 2.4. Using numerical simulations, we then demonstrate in Section 3 that our ML architecture is able to successfully recover EOS material parameters, initial shell profiles, and velocity to reproduce the observed shock and edge features, demonstrating that the time series of shock profiles does indeed encode sufficient information to infer these parameters with high accuracy. Moreover, in numerical simulation of HEDP, EOS is typically represented via underlying model assumptions or tabulated data models. We demonstrate that an expansive parameter model is sufficient to represent unknown models, in the sense that we can use an ML architecture trained on an underlying EOS model, $M_{1}$, and output parameters of $M_{1}$ for a time-series of shock profiles generated with a different underlying EOS model, $M_{2}$, that lead to a consistent time-series of shock profiles as those originally generated based on $M_{2}$. In this sense, our ML model is learning structure from the underlying physics, above a specific choice of EOS parameterization needed for numerical simulation.
@@ 2.1 Generation of density time series
As a representative problem of a double shell ICF capsule implosion, we study shock propagation in a time-dependent pseudo-3D density profile, created by the implosion of a perturbed spherical metallic shell into a medium of gas. We utilize air in lieu of D-T in order to enable examination of the out-going shock which would otherwise be obscured by the burning plasma once the hot spot is formed. We rely on the effects of a non-spherical perturbation and variation in initial velocity to provide distinct behavior in the late-time shock evolution. We assume azimuthal symmetry, therefore the density at any time can be described in 2D cylindrical coordinates $(r,z)$, but the solution remains 3-dimensional via the symmetry assumption. Additionally, we focus on the Mie-Grüneisen (MG) EOS model. We then generate a large set of data for variations in the inner shell perturbation, initial shell velocity, and EOS parameters, where each data sample contains a time series of the resulting hydrodynamic density field. Simulations are performed on a $440\times 440$ uniform Cartesian grid on a computational domain given by the quarter-plane $[0,L]\times[0,L]$, where $L=341$ $\mu$m. The uniform grid cell size is $\Delta r=\Delta z=\frac{L}{440}$. The metallic shell is made of Tantalum and its density is initially uniform at a value of 16.65 g/cc.
The shell perturbations are specified by adding harmonic perturbations to the inner surface of a spherical Tantalum shell, which can be described as the set of coordinates $(r_{\rm in}(u),z_{\rm in}(u))$ satisfying $r_{\rm in}(u)^{2}+z_{\rm in}(u)^{2}=\left(R_{\rm in}+\sum_{k=1}^{8}F_{k}\cos(2ku)\right)^{2}$, where $u\in[0,\pi/2]$, ${R_{\rm in}=248}$ $\mu$m, $F_{k}$, $k=1,\dots,8$, are coefficients of the perturbation corresponding to the $k^{\rm th}$ cosine harmonic. The outer surface of the shell is a sphere with radius ${R_{\rm out}=310}$ $\mu$m. There are 20 different inner surface perturbation profiles considered in our dataset. The corresponding coefficients are recorded in Table 4. Figure 3 presents an initial perturbation given to the interior shell. As an initial condition, the shell is given a uniform implosion velocity, $v_{\rm impl}$, in the direction of the origin to initiate an implosion. We include 4 choices of implosion velocity in our dataset.
@@
$c_{V}$ $[{\rm erg}\;{\rm g}^{-1}\;{\rm eV}^{-1}]$ $1.6\times 10^{10}$ $1.76\times 10^{10}$ $1.44\times 10^{10}$
@@ 4 Discussion
Our model problem consists of hydrodynamic simulations of an implosion of a nearly-spherical metallic shell. In our dataset, various equation of state (EOS) parameters and initial conditions on the surface and initial velocity are varied to produce multiple realizations of density field time series. For each simulation, we generated a time series of synthetic radiographs and extracted shock and edge features in the form of cosine harmonic coefficients. Through our results, we demonstrated that our approach is capable of recovering the EOS and initial condition parameters with reasonable accuracy. We also show that the parameter estimates can be successfully used in a surrogate model for the forward problem to accurately estimate the shock and edge features. Additionally, we show that the estimated parameters can be used in a (full-order) hydrodynamic solver to produce density fields with acceptable quantitative accuracy in terms of RMSE and errors in peak-to-trough distance of the RMI surface. We also investigated the ability of the model to estimate parameters when the reference simulation is generated with a different EOS model. Our findings imply that a sufficiently expansive parameter model is capable of representing unknown models.
@@ Appendix B Radiograph-to-Features Network (R2FNet)
[…]

### v2

# Physics consistent machine learning framework for inverse modeling with applications to ICF capsule implosions

@@ Abstract
In high energy density physics (HEDP) and inertial confinement fusion (ICF), predictive modeling is complicated by uncertainty in parameters that characterize various aspects of the modeled system, such as those characterizing material properties, equation of state (EOS), opacities, and initial conditions. Typically, however, these parameters are not directly observable. What is observed instead is a time sequence of radiographic projections using X-rays. In this work, we define a set of sparse hydrodynamic features derived from the outgoing shock profile and outer material edge, which can be obtained from radiographic measurements, to directly infer such parameters. Our machine learning (ML)-based methodology involves a pipeline of two architectures, a radiograph-to-features network (R2FNet) and a features-to-parameters network (F2PNet), that are trained independently and later combined to approximate a posterior distribution for the parameters from radiographs. We show that the machine learning architectures are able to accurately infer initial conditions and EOS parameters, and that the estimated parameters can be used in a hydrodynamics code to obtain density fields, shocks, and material interfaces that satisfy thermodynamic and hydrodynamic consistency. Finally, we demonstrate that features resulting from an unknown EOS model can be successfully mapped onto parameters of a chosen analytical EOS model, implying that network predictions are learning physics, with a degree of invariance to the underlying choice of EOS model. To the best of our knowledge, our framework is the first demonstration of recovering both thermodynamic and hydrodynamic consistent density fields from noisy radiographs.
@@ 1.1 Inferring physical parameters from radiographic data
Consequently, characterization of the initial conditions responsible for the instability as well as material properties characterizing the growth rates demands new techniques to solve this inverse problem. One such approach to resolve this difficulty has recently been proposed,[59] which utilizes the robust features of the outgoing shock to characterize the growth rates of the instability via a density reconstruction of the radiographic images. The success of this methodology is founded in the fact that the outgoing shock encodes sufficient information to enable machine learning techniques to learn a mapping between a sequence of outgoing shocks and the corresponding density fields. In this work we propose to utilize the outgoing shock as well as outer material edge as robust features to enable parameter estimation and estimation of the initial conditions. Using the recovered parameters in a hydrodynamic solver we then recover thermodynamically and hydrodynamically consistent density fields.
@@ 1.2 A machine learning (ML)-based inverse mapping based on shock and outer edge features
Here we develop a framework to construct machine-learning-based inverse mappings directly from experimental radiographic images to the underlying physical parameters and initial perturbations in ICF settings systems. Extraction of the parameters and initial perturbations enable subsequent use in a hydrodynamics solver to obtain density fields, shocks, and material interfaces that satisfy hydrodynamic and thermodynamic consistency to numerical accuracy. We assume a governing physical model given by the compressible Euler equations, dropping higher-order effects of radiation for ease of testing and development, analogous to recent work [60], and we assume unknown initial conditions and material properties for the governing Euler equations. As such, we build upon previous work that demonstrates the ability to learn parameters and density fields from a sequence of radiographic features for diagnosing asymmetries in the drive of ICF capsules by utilizing an inert gas as a surrogate for the D-T fuel to enable extraction of the outgoing shock [61, 47, 62]. We note that recent work at the National Ignition Facility has been performed using a silicon dopant  [63]. Furthermore, by subsequent design optimization to minimize asymmetries, improved neutron yields can be realized. Our specific problem description is one in which an ICF-like shell is imploded with an initially perturbed surface as depicted in Figure˜3. Upon collapse, the generation of a shock forms on axis and subsequently rebounds, interacting with the perturbed surface of the shell and initiates a RM instability. We posit that the outer material edge, outgoing shock, and their evolution encode sufficient information from the instability to identify the underlying simplified Mie-Grüneisenn EOS parameters and structure of the initial shell perturbation and velocity. Moreover, the shock and edge profiles are some of the few robust and identifiable features in dynamic radiographic images of HEDP experiments, which in general are subject to noise and scattering effects, and correspond to a projected areal mass rather than a primary hydrodynamic variable such as density. Thus broadly, we aim to calibrate material models and initial conditions to be consistent with the outgoing shock and edge profile, a problem of data-assimilation [64]. There are many approaches to data assimilation, and here we review some of the prominent techniques, before concluding with our specific machine-learning-based approach.
@@
The emergence of machine learning has has brought about increased attention to data-driven approaches in ICF, high energy density, and plasma physics research. Indeed, machine learning has been used to examine performance and sensitivity of ICF implosions [71]. Furthermore, Bayesian methods [72, 73, 74] have been used with relatively simple physics models to infer parameters in implosion experiments, such as laser-driven opacity and equation-of state from experiments. For an introduction and review of current work, we recommend the review paper by Knapp and Lewis [75].
@@
Consequently, in this work we develop machine learning architectures that directly take in a low-dimensional time-series of radiographic images, extract the outgoing shock and outer material edge profiles, and estimate the corresponding EOS parameters as well as initial shell conditions. An additional benefit of this approach is that the estimated parameters can then be plugged back into a hydrodynamics code to yield physically admissible density field reconstructions of the radiographic images. Such physical consistency is lacking in virtually all other reconstruction techniques currently used. The generation of data is discussed in Section˜2 and the machine learning (ML)-based parameter estimation pipeline is introduced in Section˜2.4. Using numerical simulations, we then demonstrate in Section˜3 that our ML architecture is able to successfully recover EOS material parameters, initial shell profiles, and velocity to reproduce the observed shock and edge features, demonstrating that the time series of shock profiles does indeed encode sufficient information to infer these parameters with high accuracy. These parameters are then demonstrated to yield accurate and physically admissible feature and density reconstructions by evaluating the forward hydrodynamics code using predicted parameters. Moreover, in numerical simulation of HEDP, EOS is typically represented via underlying model assumptions or tabulated data models. We demonstrate that an expansive parameter model is sufficient to represent unknown models, in the sense that we can use an ML architecture trained on an underlying EOS model, $M_{1}$, and output parameters of $M_{1}$ for a time-series of shock profiles generated with a different underlying EOS model, $M_{2}$, that lead to a consistent time-series of shock profiles as those originally generated based on $M_{2}$. In this sense, our ML model is learning structure from the underlying physics, above a specific choice of EOS parameterization needed for numerical simulation.
@@ 2.1 Generation of density time series
[…]
Refer to caption(a)Refer to captionxxzzyy(b)Refer to caption(c)Refer to caption(d)ablatorTantalum shell
@@
$c_{V}$ ${\rm erg}\;{\rm g}^{-1}\;{\rm eV}^{-1}]$ $1.6\times 10^{10}$ $1.76\times 10^{10}$ $1.44\times 10^{10}$
@@ 4 Discussion
Our model problem consists of hydrodynamic simulations of an implosion of a nearly-spherical metallic shell. In our dataset, various equation of state (EOS) parameters and initial conditions on the surface and initial velocity are varied to produce multiple realizations of density field time series. For each simulation, we generated a time series of synthetic radiographs and extracted shock and edge features in the form of cosine harmonic coefficients. Through our results, we demonstrated that our approach is capable of recovering the EOS and initial condition parameters with reasonable accuracy. We also show that the parameter estimates can be successfully used in a surrogate model for the forward problem to accurately estimate the shock and edge features. Additionally, we show that the estimated parameters can be used in a (full-order) hydrodynamic solver to produce thermodynamically and hydrodynamically consistent density fields with acceptable quantitative accuracy in terms of RMSE and errors in peak-to-trough distance of the RMI surface. To the best of our knowledge, our framework is the first to recover thermodynamically and hydrodynamically consistent density reconstruction from noisy radiographs.
We also investigated the ability of the model to estimate parameters when the reference simulation is generated with a different EOS model. Our findings imply that a sufficiently expansive parameter model is capable of representing unknown models. We note that there are other forms of model mismatch that manifest in experimental data, such as radiographic noise and scatter. While the application of our approach to experimental data is the subject of our future work, we previously demonstrated in [59] that our method of using the shock and edge to reconstruct density fields was robust in the presence of out-of-distribution noise and scatter. More recently, a comprehensive investigation was performed to examine the impact of out-of-distribution noise and scatter [76]. The results again confirmed that the sparse features are robust and consequently accurate density fields can be obtained from the dynamic sequence of these features with significant out-of-distribution scatter and noise.
Since the F2PNet is generative, it is a useful tool in the uncertainty analysis of predicted parameters. Inspired by the conditional variational autoencoder (cVAE), the F2PNet architecture contains a decoder which applies a nonlinear transformation to features and a sample of multi-dimensional Gaussian noise to obtain a predicted distribution for the parameters. While the decoder is capable of producing multi-modal distributions, we largely observe that F2PNet trained on the dataset considered in this paper predicts uni-modal distributions for the parameters. Typically, the characteristics of the predicted parameter distribution depend on the weighting of the KL divergence loss term and dimensionality of the noise in the latent space. However in our investigations, we observed insensitivity with respect to the choice of these meta parameters and we therefore speculate that the observed uni-modal predictions are due to the data providing sufficient information content to closely identify the sensitive hydrodynamic parameters. Our speculation is also supported by our past work [47], where we investigated recovering the density field in the vicinity of the Richtmeyer-Meshkov instability from the same hydrodynamic features dataset considered in this paper using a generative cVAE-based vision transformer. In this work, we observed that sampling the latent space of the generative network for a given feature produced a distribution of quantitatively close density fields with topologically similar Richtmeyer-Meshkov instabilities.
@@ Appendix B Radiograph-to-Features Network (R2FNet)
The radiograph-to-features network (R2FNet) is illustrated in Figure 12 and is built entirely from convolutions and fully connected linear layers. The first component of the architecture is a convolutional encoder which compresses the series of radiographs from their large ambient dimension to a much smaller latent dimension. The downsampling blocks used in this architecture are identical to the residual blocks used in [83], and are very similar to those found in BigGAN [84]. The network input consists of the four radiographs stacked together so that each timestep is represented by one input channel. The first downsampling block then increases the channel dimension to 64. Each further downsampling block adds an additional 64 channels, and all the downsampling blocks reduce the spatial dimensions by a factor of 2. In the downsampling blocks, the first ${\times}3$ convolution increases the number of channels. Both networks use 7 downsampling blocks. The outputs of the encoder are then flattened into a vector. Subsequently, there are 3 linear layers - the first of which reduces the dimension by a factor of 16, the second maintains this dimension, and the final layer outputs the parameter predictions.
@@ Appendix C Features-to-Parameters Network (F2PNet)
Data availability
The datasets analyzed during the current study are available from the corresponding author on reasonable request.